a pudding recipe for 50 people calls for 4 cups of sugar. each bag of sugar contains 6 cups. how many bags of sugar will be needed to make this recipe for 300 people?
Answer #1
300/50 = 6

6 * 4 = 24 cups needed

24/6 = 4 bags needed
answered by: Asia K.